OpenStudy (anonymous):

Solve: -1/2 x + 1 = -x + 8

OpenStudy (anonymous):

Choices: 7 1/2 14 18

OpenStudy (anonymous):

Ok, so first you can cancel out any numbers you don't need. \[-\frac{ 1 }{ 2 }x + 1 -1 = -x + 8 - 1\] Which leaves you with \[-\frac{ 1 }{ 2 }x = -x - 7\]

alones (alones):

Let's solve your equation step-by-step. −1/2x+1=−x+8 Step 1: Add x to both sides. −1/2x+1+x=−x+8+x 1/2x+1=8 Step 2: Subtract 1 from both sides. 1/2x+1−1=8−1 1/2x=7 Step 3: Multiply both sides by 2. 2*(1/2x)=(2)*(7) x=14 Answer: x=14 Hope it helps =]

OpenStudy (anonymous):

Thank you guys so much :3

OpenStudy (anonymous):

obviously you have a quicker mind then me AloneS haha

OpenStudy (anonymous):

And you're welcome :)

alones (alones):

Ha ha some day you'll have it too<33 @JJAwesome101 and you'll type faster :3

OpenStudy (anonymous):